[Hallc_running] [New Logentry] RC Update

hszumila at jlab.org hszumila at jlab.org
Tue Nov 9 08:45:01 EST 2021

Logentry Text:
We completed the production running yesterday. 

There was an incident yesterday at 4pm where 70uA was momentarily put on the Al dummy target. This caused a down time last night while MCC investigated the procedure. 
Also, we had an issue talking to the SHMS dipole this morning that also caused a downtime. Thanks to Ellen for making the access and getting the PLC back on comm. 

Currently, we are now taking the elastic singles. I anticipate this will go on through the end of swing. After this we should be able to take a comparison Al dummy run to check the charge normalized yield before proceeding to the delta scan. 

The current plan is to have beam removed tomorrow at 8:30am to begin the pass change. Radcon will come to sweep our Hall for Controlled Access at that time. We will likely be down for most of the day shift for an aerogel swap. 

Today's summary and plan: https://hallcweb.jlab.org/wiki/index.php/RC_Meeting,_Tuesday,_November_9,_2021

Connection details for 10am meeting: https://bluejeans.com/861439920


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3939261
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20211109/77980f01/attachment.html>

More information about the Hallc_running mailing list